The metabolic rate R ( kJ day-1) of mammals varies with their body

mass M (kg) according to the power law R=k M3/4. where k is a constant.

 

A common shrew has a mass of 9 g and a metabolic rate of 7 kJ day-1.

 

(a) What is the metabolic rate in kJ day-1 of a dog whose mass is 5 kg? Give the answer to at least 3 significant figures.

 

(b) What is the mass in kg of a cat with a metabolic rate 560 kJ day-1? Give the answer to at least 3 significant figures.


1
Expert's answer
2021-11-11T02:47:01-0500

a)Metabolic rate in kJ day-1 of a dog whose mass is 5 kg;


"=\frac{5kg\times 7KJday^{-1}}{9\times10^{-3}Kg}"


=3,888.889 KJ day-1


b)mass in kg of a cat with a metabolic rate 560 kJ day-1


"=\frac{560KJday^{-1}\times 9\times10^{-3}Kg}{7KJday^{-1}}"


=0.720 Kg